# Correcting Annotator Bias in Training Data: Population-Aligned Instance Replication (PAIR)

### v1

@@ A.2 Model Training Details
A.3 Additional Results
We present the results of F1 scores for filtered tweets ($0.4\leq p_{i}\leq 0.6$) in Figure 6.
@@ A.3 Additional Results
Figure 6: Model F1 scores for filtered tweets ($0.4\leq p_{i}\leq 0.6$), by dataset and bias ($\beta$)
A.4 Additional Metrics
We also use the Expected Calibration Error (ECE) metric to evaluate model calibration. This metric quantifies calibration quality by measuring the weighted average absolute difference between accuracy and predicted confidence across $M$ bins:
$\text{ECE}=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\cdot\left|\text{acc}(B_{m})-\text{conf}(B_{m})\right|$ (4)
where $B_{m}$ is the set of samples in the $m$-th bin, $|B_{m}|$ its size, $n$ the total number of samples, $\text{acc}(B_{m})$ the accuracy, and $\text{conf}(B_{m})$ the mean predicted confidence in $B_{m}$.
Figure 7 shows the results. Consistent with our main findings in §4, we observe that for OL labels, the ECE values are relatively stable on the Balanced and Adjusted datasets ($t=-0.48,p=0.64$), with divergence with $\beta=0.15$. This contrasts with the significant divergence in ECE values for the Unbalanced subsets as the $\beta$ value increases.
@@ A.4 Additional Metrics
Figure 7: Model ECE metrics, by dataset and bias ($\beta$)
Additionally, for model performance, we use another metric ROC-AUC in Figure 8. The results are very similar to the F1 results in §4.
Figure 8: Model ROC-AUC metrics, by dataset and bias ($\beta$)
A.5 Adjustment Details
The adjustment we use to make the labeler pool more representative of the target population is a form of pseudo-population generation (Quatember, 2015). We create the pseudo-population by first constructing post-stratification weights, performing weight normalization to ensure the sum of the weights equals the size of the target population, and then duplicating each observation proportionally to its weight via deterministic replication.
Post-stratification Bethlehem et al. (2011b) is a method of statistical adjustment that makes a selected sample more closely resemble a target population. Post-stratification requires population level totals or proportions and corresponding case-level observations of the same characteristic. to be available for each sample stratum that will be used in the weighting. The weight for unit $i$ is determined for each stratum by:
$w_{i}=\frac{P_{s}}{S_{s}}$ (5)
where $P_{s}$ is the true population proportion (or total) for stratum (or group) $s$ and $S_{s}$ is the sample proportion (or total) for stratum $s$. In our case, the strata of interest was a single variable (labeler Types A & B). However, post-stratification can involve multiple variables if their joint distribution is known at the population level.
where $T$ is the target total. Since we want the Adjusted dataset to match the size of our gold standard Balanced dataset, the target total for the simulation was 12 labels per tweet.
Lastly, to construct a pseudo-population from our weighted data, we perform deterministic replication by replicating each unit $n_{i}$ times. In this initial work, we appreciated the simple interpretation and reproducibility of this approach. However, researcher may prefer other approaches, such as replication via resampling, if they are interested in how the adjustment varies across samples.
In our case, after post-stratification and weight normalization, each Type A label in the Unbalanced 1 dataset receives a weight of 1 and each Type B label receives a weight of 2. This resulted in an Adjusted dataset where each of the Type A labels stays the same and each of the Type B labels is duplicated once.

### v2

@@ Appendix B Model Training Details
Appendix C Ablation Study 1: HS Results
We show the results on HS labels in this section. We construct the HS probabilities for the A and B annotators are defined in the same way as the OL probabilities: $p_{i,HS}^{A}=\text{max}(p_{i,HS}-\beta,0)$, $p_{i,HS}^{B}=\text{min}(p_{i,HS}+\beta,1)$. We also construct the four data sets (Representative, Non-representative 1, Non-representative 2, Adjusted) in the same way we did for the OL case.
Label Distribution during Simulations.
Figure 4 shows the percentage of tweets labeled HS in the four datasets for each value of $\beta$. As expected, the percentage in the Adjusted dataset is similar to that in the Representative dataset for all values of $\beta$, and the two unbalanced datasets have lower rates of HS, because they overrepresent the A annotators, who are less likely to label HS.
Figure 4: Percentage of tweets annotated as HS, by dataset and bias ($\beta$)
HS is rare in our dataset (16.7% of HS labels are yes, Table 1). And, our simulation strategy overrepresents $A$ annotators in the two Non-representative data sets, who are less likely to perceive HS (Table 2). For these reasons, a higher proportion of $p_{i,HS}^{A}$ are 00 while the $p_{i,HS}^{B}$ probabilities continue to increase. This issue leads the proportion of “yes” labels in the Representative and Adjusted datasets to increase with $\beta$ in the HS data set, which have more B labels than the unadjusted datasets.
Results.
Figure 5 (1) contains the ACB results and Figure 5 (2) the F1 score results for the HS datasets. The PAIR approach does not improve calibration or accuracy: the adjusted model performs similarly to the Non-representative models. This effect is likely due to the combination of label rarity and our simulation design. With few positive labels to begin with, the impact of varying annotator characteristics through our $\beta$ parameter may be overwhelmed by the baseline scarcity of hate speech annotations. Ablation Study 2 (Appendix §D) addresses this point.
Figure 5: Metrics for HS models, by dataset and bias ($\beta$)
Appendix D Ablation Study 2: Difficult Tweets
Our simulations assumed that all tweets are impacted the same way (Eq. (2)), which is an oversimplification. More realistically, annotator characteristics likely have more impact for ambiguous tweets. For this reason, we repeat model training and recompute metrics for those tweets where $0.4\leq p_{i}\leq 0.6$. This approach not only focuses on those tweets where annotator characteristics likely play a larger role, it also eliminates the floor and ceiling effects in Eq. 2). The filtered datasets contain 267 (OL) and 360 (HS) tweets.
Figures 6 and 7 show the results for two metrics (ACB, F1) for filtered OL and HS labels. In Figure 6(1): the Representative and Adjusted models have similar ACB and are lower than the Non-representative models. The F1 scores do not show differences between the models. These results are similar to those on the full set of tweets (Figure 3). In the two HS figures (7), we see signs that the Representative and Adjusted models perform similarly, and better than the two Non-representative models, on both metrics. These results are more promising than those on the full set of tweets (Figure 5).
@@ Appendix D Ablation Study 2: Difficult Tweets
Figure 6: Metrics for OL models for filtered tweets ($0.4\leq p_{i,OL}\leq 0.6$), by dataset and bias ($\beta$)
@@
Figure 7: Metrics for HS models for filtered tweets ($0.4\leq p_{i,HS}\leq 0.6$), by dataset and bias ($\beta$)